### v2

# Eigenvalues of the Neumann magnetic Laplacian in the unit disk

Abstract
In this paper, we study the first eigenvalue of the magnetic Laplacian with Neumann boundary conditions in the unit disk $\mathbb{D}$ in $\mathbb{R}^{2}$. There is a rather complete asymptotic analysis when the constant magnetic field tends to +$\infty$+ ∞ and some inequalities seem to hold for any value of this magnetic field, leading to rather simple conjectures. Our goal is to explore these questions by revisiting a classical picture of the physicist D. Saint-James theoretically and numerically. On the way, we revisit the asymptotic analysis in light of the asymptotics obtained by Fournais-Helffer, that we can improve by combining them with a formula stated by Saint-James.
1 Introduction
In this paper, we study the first eigenvalue of the magnetic Laplacian with Neumann boundary conditions in the unit disk $\mathbb{D}$ in $\mathbb{R}^{2}$. Let us formulate the problem more precisely. Using the standard choice
$\mathbf{A}:x=(x_{1},x_{2})\in\mathbb{R}^{2}\mapsto(-x_{2}/2,x_{1}/2)\in\mathbb{R}^{2}$
for the so-called magnetic potential, we define the self-adjoint operator $H_{\beta}$ (with $\beta\in\mathbb{R}$) by
${H_{\beta}u}:=(i\nabla+\beta\,\mathbf{A}(x))^{2}u=-\Delta u+2i\,\beta\,\mathbf{A}\cdot\nabla u+\beta^{2}|\mathbf{A}|^{2}u\,,$
for all $u$ in the domain
$\mathcal{D}:=\{u\in H^{2}(\mathbb{D})\,:\,\nabla u\cdot\vec{\nu}=0\mbox{ on }\partial\mathbb{D}\},$
where $\vec{\nu}$ is the outward-pointing unit normal. It is well known that $H_{\beta}$ is a self-adjoint operator, non-negative (positive if $\beta\neq 0$), with compact resolvent. Let $\lambda(\beta)$ denote its lowest eigenvalue. Our goal is to understand in detail the dependence of $\lambda(\beta)$ on the parameter $\beta$ (interpreted as the strength of the magnetic field).
The analysis of this Neumann problem is strongly motivated by the analysis of surface superconductivity in physics and the Neumann boundary condition is crucial (see [29]). At the mathematical level the case of the disk is a way towards understanding the role of the curvature of the boundary. We refer to [3], [20] and to [15] for a presentation of the state of the art in 2014 including (to mention some of them) works by Bernoff-Sternberg [3], Bauman-Phillips-Tang [1] (where the analysis on the disk plays an important role) and Lu-Pan [25].
Notice that the Dirichlet problem has been analyzed intensively in the last years with similar techniques (see [30, 2] and references therein) but let us emphasize that the phenomena observed in the Neumann case are quite different.
For the analysis of the spectrum $H_{\beta}$, using the radial symmetry of the domain $\mathbb{D}$, we pass to the polar coordinates
$(x_{1},x_{2})=(r\cos\theta,r\sin\theta)$
and use a decomposition in Fourier series with respect to the variable $\theta$. We are led to consider the family of $1D$-operators indexed by $n\in\mathbb{Z}$
$H_{n,\beta}:=-\frac{d^{2}}{dr^{2}}-\frac{1}{r}\frac{d}{dr}+\left(\frac{\beta}{2}r-\frac{n}{r}\right)^{2}\,,$ (1.1)
[…]
For each of these operators, we look at the lowest eigenvalue $\lambda(n,\beta)$ as a function of $\beta$. It follows from standard perturbation theory of self-adjoint operators that each map $\beta\mapsto\lambda(n,\beta)$ is real-analytic. We then get the lowest eigenvalue from
$\lambda(\beta)=\inf_{n\in\mathbb{Z}}\lambda(n,\beta)\,.$ (1.2)
As explained in Section 2, we can restrict ourselves to $\beta\geq 0$ and $n\in\mathbb{N}$ without loss of generality.
We often find it more convenient to consider auxiliary maps, defined in $(0,+\infty)$ by
$\beta\mapsto\eta(n,\beta):=\frac{\lambda(n,\beta)}{\beta}$
and
$\beta\mapsto\eta(\beta):=\frac{\lambda(\beta)}{\beta}.$
Our first goal (suggested by previous work and by numerical computations) is to understand the intersection of the curves $\beta\mapsto\lambda(n,\beta)$ and $\beta\mapsto\lambda(n+1,\beta)$ (or, equivalently, of the curves $\beta\mapsto\eta(n,\beta)$ and $\beta\mapsto\eta(n+1,\beta)$). We give a complete proof of a formula due to the physicist D. Saint-James [28].
Theorem 1.1.
Let $n\in\mathbb{N}$ and let us assume that $\beta^{*}$ is a solution in $[0,+\infty)$ of the equation
$\lambda(n,\beta)=\lambda(n+1,\beta).$
Then, with the notation
$\eta^{*}:=\eta(n,\beta^{*})=\eta(n+1,\beta^{*}),$
we have
$\beta^{*}=2\eta^{*}+2n+1+\sqrt{(2\eta^{*}+1)^{2}+8n\eta^{*}},$
and, as a consequence, $\beta^{*}>2(n+1)$.
We then try to understand which $n$ realizes the infimum (1.2) according to the value of $\beta$. We combine the results of M. Dauge and the first author [9, 10], on eigenvalues variation for Sturm-Liouville operators, with Theorem 1.1, to obtain the following result.
Theorem 1.2.
For any $n\in\mathbb{N}$, there exists a unique solution $\beta_{n}$ in $[0,+\infty)$ of the equation
$\lambda(n,\beta)=\lambda(n+1,\beta).$
In addition, the sequence $(\beta_{n})_{n\in\mathbb{N}}$ is strictly increasing, and
  1. (i)
    for all $\beta\in[0,\beta_{0}]$, $\lambda(\beta)=\lambda(0,\beta)$;
  2. (ii)
    for all $n\in\mathbb{N}\setminus\{0\}$, for all $\beta\in[\beta_{n-1},\beta_{n}]$, $\lambda(\beta)=\lambda(n,\beta)$.
Furthermore, the equation $\lambda(k,\beta)=\lambda(\beta)$, with $\beta\geq 0$, is satisfied only for $k=n$ when $\beta\in(\beta_{n-1},\beta_{n})$, with $n>0$, (only for $k=0$ when $\beta\in[0,\beta_{0}^{*})$) and only for $k\in\{n,n+1\}$ when $\beta=\beta_{n}$
Finally, we study the following conjectures.
Conjecture 1.3.
For all $\beta>0$,
$\eta(\beta)<\Theta_{0}\,.$
In the above statement, $\Theta_{0}$ is a well-known universal constant, called the De Gennes constant since it was introduced by him [8, 29] in the sixties in the context of Surface Superconductivity (see for mathematical studies [4], and later the book [15] and references therein). This constant is defined
$\Theta_{0}:=\inf_{\xi\in\mathbb{R}}\lambda^{DG}(\xi)\,,$ (1.3)
where $\lambda^{DG}(\xi)$ is the lowest eigenvalue of the Neumann realization of the harmonic oscillator
$\mathfrak{h}_{0}(\xi):=D_{t}^{2}+(t+\xi)^{2}\mbox{ on }\mathbb{R}^{+}\,.$ (1.4)
Numerically [^4] See [27]., $\Theta_{0}\simeq 0.590106$ and its first role is to give the asymptotics
$\lambda(\beta)=\Theta_{0}\beta+o(\beta)$ (1.5)
as $\beta$ tends to +$\infty$+ ∞, so that $\eta(\beta)\to\Theta_{0}$.
This property has been known since the nineties and we actually have more accurate asymptotics for the disk, playing an important role in the analysis of more general domains. Our main interest is in the inequality, which is equivalent to $\lambda(\beta)<\Theta_{0}\,\beta$ and is an improvement of the result $\lambda(\beta)<\beta$, proved in [6, Theorem 2.1] for a general domain in $\mathbb{R}^{2}$.
We show in Section 3.5 that Conjecture 1.3 can be deduced from the following other conjecture.
Conjecture 1.4.
The sequence $(\eta_{n}^{*})_{n\in\mathbb{N}}$ is strictly increasing, where
$\eta_{n}^{*}:=\eta(n,\beta_{n})=\eta(n+1,\beta_{n})$
and $\beta_{n}$ is defined in Theorem 1.2.
We also investigate the so-called strong diamagnetism in $\mathbb{D}$, which states as a conjecture:
Conjecture 1.5.
The map $\beta\mapsto\lambda(\beta)$ is monotonically increasing in the interval [0,+[[0,$\infty$[[ 0 , + ∞ [.
It follows from the result of S. Fournais and B. Helffer for a general domain in $\mathbb{R}^{2}$ [14] that the map is ultimately increasing, that is, there exists some $\overline{\beta}>0$ such that $\beta\mapsto\lambda(\beta)$ is increasing in $[\overline{\beta},+\infty[$. The point is to show that, in the case of the disk, the map is increasing in the whole interval [0,+[[0,$\infty$[[ 0 , + ∞ [. Let us note that this property is known to hold when the Neumann boundary condition is replaced with Dirichlet (see the PhD thesis of S. Son [30, Theorem 3.3.4]).
In Sections 6 and 4, we present asymptotic and numerical computations supporting these conjectures. We also show how the combination of the results in [14] with the approach developed in the first sections leads to new relations and permits to control the validity of the conjectures in the large magnetic field limit.
Acknowledgements.
This problem has been discussed along the years with many colleagues sharing their ideas. Many thanks to B. Colbois, S. Fournais, A. Kachmar, G. Miranda, M. Persson-Sundqvist, L. Provenzano, and A. Savo for their contributions.
C. Léna received financial support from the INdAM GNAMPA Project “Operatori differenziali e integrali in geometria spettrale”.
2 Eigenvalue curves
We have, for the operator $H_{n,\beta}$ defined by Equation (1.1),
$H_{n,\beta}=H_{-n,-\beta},$
so that, from Equation (1.2), we get
$\lambda(-\beta)=\lambda(\beta)\,.$
In addition, a direct computation shows that
$H_{-n,\beta}=H_{n,\beta}+2n\beta,$
and therefore, in particular
$\lambda(-n,\beta)=\lambda(n,\beta)+2n\beta.$
For $\beta>0$, this implies that the infimum
$\lambda(\beta)=\inf_{n\in\mathbb{Z}}\lambda(n,\beta)$
cannot be realized by a negative integer. Thus, if we want to study the map $\beta\mapsto\lambda(\beta)$, it is enough to study the maps $\beta\mapsto\lambda(n,\beta)$ for $n\geq 0$ and $\beta\geq 0$.
For $\beta=0$, we recover the usual Laplacian, with Neumann boundary conditions. We therefore assume in the rest of this section that $\beta>0$. The Sturm-Liouville problem corresponding to $H_{n,\beta}$ is
$\left\{\begin{array}[]{rl}-f^{\prime\prime}-\frac{1}{r}f^{\prime}+\left(\frac{n}{r}-\frac{\beta}{2}r\right)^{2}f=&\lambda f\mbox{\ \ in \ \ }(0,1)\,,\\ f(0)=0\mbox{ if }n>0,&f^{\prime}(0)=0\mbox{ if }n=0\,,\\ f^{\prime}(1)=&0\,.\end{array}\right.$ (2.1)
We write $r\mapsto f_{n,\beta}(r)$ for the positive and normalized eigenfunction of (2.1) associated with $\lambda(n,\beta)$.
We can express it using the Kummer function $z\mapsto M(a,b,z)$ (with the conventions and notation of the Digital Library of Mathematical Functions [11, Chapter 13]):
$f_{n,\beta}(r)=C_{n}^{\beta}r^{n}e^{-\frac{\beta}{4}r^{2}}M\left(\frac{1}{2}-\frac{\lambda}{2\beta},n+1,\frac{\beta}{2}r^{2}\right)$ (2.2)
where $\lambda=\lambda(n,\beta)$.
Notice (cf [26]) that when $a<c$
[…]
Alternatively, we can use the Whittaker function $z\mapsto M_{\kappa,\mu}(z)$ (see for instance [23, Equation (A.5)] [^5] The authors thank Mikael Persson Sundqvist for communicating to them unpublished notes and programs.):
$f_{n,\beta}(r)=\frac{D_{n}^{\beta}}{r}M_{\frac{\lambda}{2\beta}+\frac{n}{2},\frac{n}{2}}\left(\frac{\beta}{2}r^{2}\right).$ (2.4)
The constants $C_{n}^{\beta}$ and $D_{n}^{\beta}$, depending on $n$ and $\beta$, are chosen so that
[…]
The eigenvalues of (2.1) (and in particular the first eigenvalue, $\lambda(n,\beta)$) are the roots of the equation
$(f_{n,\beta})^{\prime}(1)=0$ (2.5)
(where the left-hand side is understood as the function of $\lambda$ defined by Equation (2.4)). As an illustration, we present in Figure 1 the curves
$\beta\mapsto\eta(n,\beta)=\frac{\lambda(n,\beta)}{\beta}$
for $n=0,1,\dots,20\,,$ obtained by solving Equation (2.5) numerically (with computations done in Mathematica [^6] Wolfram 14.0 For Desktop. Version: 14.0.0.0.).
Figure 1: Plot of $\beta\mapsto\eta(n,\beta)=\frac{\lambda(n,\beta)}{\beta}$ for $n$ from 0 to 20. The black dotted line and the black dashed line are the constants $1$ and $\Theta_{0}$ respectively.
In order to study the functions $\beta\mapsto\eta(n,\beta)$, let us give an alternative interpretation of $\eta(n,\beta)$. We perform the change of variable $\rho=\beta^{1/2}r$ and divide by $\beta$. We obtain a new Sturm-Liouville operator $G_{n,\beta}$, which is the Dirichlet-Neumann (Neumann-Neumann if $n=0$) realization in $(0,\beta^{1/2})$ of the differential operator
$-\frac{d^{2}}{d\rho^{2}}-\frac{1}{\rho}\frac{d}{d\rho}+\Big{(}\frac{n}{\rho}-\frac{\rho}{2}\Big{)}^{2}\,.$
This new operator is self-adjoint in the Hilbert space $L^{2}((0,\beta^{1/2}),\rho\,d\rho)$ (we recall that the measure has a weight) and $\eta(n,\beta)$ is its first eigenvalue.
3 Properties of the eigenvalue curves
3.1 Asymptotic behavior
Proposition 3.1.
Under the assumption $n>0$, we have, as $\beta\to 0^{+}$,
$\eta(n,\beta)=\frac{(j_{n,1}^{\prime})^{2}}{\beta}(1+o(1)),$
where $j_{n,1}^{\prime}$ is the smallest positive zero of $J_{n}^{\prime}$, the derivative of the Bessel function of the first kind $J_{n}$.
In addition, $\eta(0,\beta)\to 0$ as $\beta\to 0^{+}$.
Proof.
By definition, $\eta(n,\beta)=\frac{\lambda(n,\beta)}{\beta}$. From standard perturbation theory, the eigenvalues of $H_{n,\beta}$ converge to the eigenvalues of $H_{n,0}$ as $\beta\to 0$. In particular, $\lambda(n,\beta)\to\lambda(n,0)$. It is easily checked that $\lambda(n,0)=(j_{n,1}^{\prime})^{2}$ (see [22] and references therein).
The last limit follows from the well-known bound $\lambda(0,\beta)\leq\frac{1}{8}\beta^{2}$. This can be obtained by using the constant function $f(r)=1$ in the associated Rayleigh quotient:
[…]
The following result is a special case of a theorem proved by A. Kachmar and G. Miranda [24, Theorem 1] with techniques inspired by [4].
Proposition 3.2.
As $\beta\to+\infty$,
$\lambda(n,\beta)=\beta-e^{-\frac{\beta}{2}}\left(\frac{\beta^{n+2}}{2^{n}n!}+o\left(\beta^{n+2}\right)\right).$
From this proposition we have, as $\beta\to+\infty$,
$\eta(n,\beta)=1-e^{-\frac{\beta}{2}}\left(\frac{\beta^{n+1}}{2^{n}n!}+o\left(\beta^{n+1}\right)\right).$ (3.1)
Notice that the Dirichlet case is considered in [2].
3.2 Variations
In Section 2, we have defined the operator $G_{n,\beta}$ and the map $\beta\mapsto\eta(n,\beta)$ for $\beta\in(0,+\infty)$ and $n\in\mathbb{N}$. However, these definitions make sense for any real $n\in[0,+\infty)$. We will use the extended definitions at some points of our analysis. In particular, it follows from standard perturbation theory for self-adjoint operators that the map $(n,\beta)\mapsto\eta(n,\beta)$ is smooth, even real-analytic, and the Feynman-Hellmann formula gives us the derivative of $n\mapsto\eta(n,\beta)$, for a given $\beta>0$:
[…]
where $g_{n,\beta}$ is the positive and normalized eigenfunction of the operator $G_{n,\beta}$ associated with $\eta(n,\beta)$.
The work of M. Dauge and the first author [9, 10] gives a rather complete picture of the variation of $\beta\mapsto\eta(n,\beta)$ in $(0,+\infty)$, for a given $n\in\mathbb{N}$ (the results actually hold without assuming that $n$ is an integer). The derivatives in the following statements are taken with respect to $\beta$.
Proposition 3.3.
Let $n\in\mathbb{N}$.
  1. (i)
    For all $\beta>0$,
    ${\eta^{\prime}(n,\beta)}=\frac{1}{2}\,\big{(}g_{n,\beta}(\beta^{1/2})\big{)}^{2}\left(\left(\frac{n}{\beta^{1/2}}-\frac{\beta^{1/2}}{2}\right)^{2}-\eta(n,\beta)\right).$
  2. (ii)
    The map $\beta\mapsto\eta(0,\beta)$ is increasing in $[0,+\infty)$.
  3. (iii)
    If $n>0$, the map $\beta\mapsto\eta(n,\beta)$ has a unique point of minimum, which is denoted by $\beta_{\min}(n)$. In addition $\beta_{\min}(n)$ is non-degenerate and satisfies
    $\beta_{\min}(n)>2n\,.$
To unify the notation, we set
$\beta_{\min}(0):=0\,.$ (3.3)
Proof.
Setting $a:=\beta^{1/2}$, $\eta(n,\beta)$ is the first eigenvalue of a Sturm-Liouville operator of the form
$-\frac{1}{\rho}\,\frac{d}{d\rho}\left(\rho\,\frac{d}{d\rho}\right)+q_{n}(\rho),$
in the weighted $L^{2}$-space $L^{2}((0,a),\rho\,d\rho)$, with Neumann boundary condition at $\rho=a$ and either Dirichlet (when $n>0$) or Neumann (when $n=0$) boundary condition at $\rho=0$. We write $\mu_{n}(a)=\eta(n,a^{2})$. As in the reference [9], the right endpoint $a$ is variable while the potential
$q_{n}(\rho)=\left(\frac{n}{\rho}-\frac{\rho}{2}\right)^{2}$
is independent of $a$.
According to a variant of the Feynman-Hellmann formula [9, Equation ($\rm{F}_{\omega}$) p. 248], for all $a>0$,
$\mu^{\prime}_{n}(a)=a\,g(a)^{2}\left(q_{n}(a)-\mu_{n}(a)\right),$
where $g=g_{n,a^{2}}=g_{n,\beta}$. Substituting for $\mu_{n}^{\prime}(\beta^{1/2})$ in
${\eta^{\prime}(n,\beta)}=\frac{d}{d\beta}\left(\mu_{n}(\beta^{1/2})\right)=\frac{1}{2\beta^{1/2}}\mu_{n}^{\prime}(\beta^{1/2}),$
we obtain Point (i). Points (ii) and (iii) are direct applications of [9, Theorem 4.2] and [9, Theorem 4.3], respectively.
As observed in Fournais–Persson-Sundqvist [19], a variant of the first Feynman-Hellmann formula (3.2) gives
[…]
and therefore:
Lemma 3.4.
If $\beta<2n$, then $\lambda^{\prime}(n,\beta)<0$.
It is also noted in [19] that, using the trial function $f(r)=r^{n}\exp(-\beta r^{2}/4)$ in the Rayleigh quotient, we get:
Lemma 3.5.
If $\beta>2n$, $\lambda(n,\beta)<\beta$
3.3 Intersections
Let us prove a useful intermediate result towards the proof of Theorems 1.1 and 1.2.
Proposition 3.6.
Let $n\in\mathbb{N}$. There exists at least one positive $\beta^{*}$ such that $\eta(n,\beta^{*})=\eta(n+1,\beta^{*})$. In addition, any $\beta$ such that $\eta(n,\beta)=\eta(n+1,\beta)$ necessarily satisfies $\beta>2n$.
Proof.
Propositions 3.1 and 3.2 imply, respectively, that $\eta(n,\beta)<\eta(n+1,\beta)$ for $\beta>0$ small enough and $\eta(n,\beta)>\eta(n+1,\beta)$ for $\beta$ large enough. By continuity, there exists at least one $\beta^{*}>0$ such that $\eta(n,\beta^{*})=\eta(n+1,\beta^{*})$.
We prove the second part of the statement by contraposition. Let us use the extended definition of $(m,\beta)\mapsto\eta(m,\beta)$ where $m$ is not necessarily an integer. If $\beta>0$ and $m>0$ are such that $\beta\leq 2m$, Formula (3.2) shows that $\partial_{m}\eta(m,\beta)>0$. In particular, if $0<\beta\leq 2n$, the map $m\mapsto\eta(m,\beta)$ is strictly increasing in $[n,n+1]$, and therefore $\eta(n,\beta)<\eta(n+1,\beta)$.
It is not clear at this stage that there is no more than one point of intersection between the two curves. We show this in the next two sections.
3.4 Saint-James Formula
Figure 2: Saint-James picture recovered by M. Persson-Sundqvist
The goal of this section is to prove Theorem 1.1. We use a representation for the first eigenfunction of $G_{n,\beta}$ deduced from (2.2) by the relation
$g_{n,\beta}(\rho)=\beta^{-1/2}f_{n,\beta}(\beta^{-1/2}\rho)\,,$
that is
$g_{n,\beta}(\rho)=K_{n}^{\beta}\rho^{n}e^{-\frac{\rho^{2}}{4}}M\left(\frac{1}{2}-\frac{\eta}{2},n+1,\frac{\rho^{2}}{2}\right)$ (3.4)
(with $\eta=\eta(n,\beta)$ and $K_{n}^{\beta}=\beta^{-(n+1)/2}C_{n}^{\beta}$, where $C_{n}^{\beta}$ is the normalization constant in Equation (2.2)).
We obtain
$(g_{n,\beta})^{\prime}(\rho)=\left(\frac{n}{\rho}-\frac{\rho}{2}\right)g_{n,\beta}(\rho)+K_{n}^{\beta}\,\rho^{n+1}e^{-\frac{\rho^{2}}{4}}\partial_{z}M\left(\frac{1}{2}-\frac{\eta}{2},n+1,\frac{\rho^{2}}{2}\right).$
Recalling the general formula for the derivative of Kummer functions [11, (13.3.4)]:
$\partial_{z}M(a,b,z)=\frac{a}{b}M(a+1,b+1,z),$
we finally get
$(g_{n,\beta})^{\prime}(\rho)=\left(\frac{n}{\rho}-\frac{\rho}{2}\right)g_{n,\beta}(\rho)+K_{n}^{\beta}\,\frac{\frac{1}{2}-\frac{\eta}{2}}{n+1}\rho^{n+1}e^{-\frac{\beta\rho^{2}}{4}}M\left(\frac{3}{2}-\frac{\eta}{2},n+2,\frac{\rho^{2}}{2}\right)$
The Neumann condition at $\beta^{1/2}$ (i.e. $(g_{n,\beta})^{\prime}(\beta^{1/2})=0$) leads, after simplifications, to the implicit equation for the eigenvalues $\eta$:
$(n+1)(n-x)M(\nu,n+1,x)+2x\nu M(\nu+1,n+2,x)=0,$
with the notation
$\nu$ $:=\frac{1}{2}(1-\eta),$
$x$ $:=\frac{\beta}{2}.$
Let $(\beta,\eta)$ be a point of intersection of the analytic curves $\beta\mapsto\eta$ for the first eigenvalue of $G_{n,\beta}$ and $G_{n+1,\beta}$. We first note that, according to Proposition 3.6, we have $\beta>2n$. As a consequence, according to Lemma 3.5, $\lambda(n,\beta)<\beta$, so that $10<\eta<$ 1. The corresponding $(x,\nu)$ satisfies the system of equations
$(n+1)(n-x)M(\nu,n+1,x)+2x\nu M(\nu+1,n+2,x)$ $=0\,,$ (3.5)
$(n+2)(n+1-x)M(\nu,n+2,x)+2x\nu M(\nu+1,n+3,x)$ $=0\,$ (3.6)
(we have $\nu=\frac{1}{2}(1-\eta)\in(0,\frac{1}{2})$, so the Kummer functions are well-defined).
Remark 3.7.
Since $(g_{n,\beta})^{\prime}(\beta^{1/2})=0$ and since $g_{n,\beta}$ is a non-zero solution of a second-order differential equation, we necessarily have $g_{n,\beta}(\beta^{1/2})\neq 0$, or equivalently $M(\nu,n+1,x)\neq 0$.
We proceed in the following way to obtain an algebraic equation satisfied by $x$ and $\nu$.
  • We use the known recursion relations for the functions $M(a,b,z)$ (see for instance [11, Section 13.3]) to express the left-hand side in the system as a linear combination of
    $X$ $:=M(\nu,n+1,x),$
    $Y$ $:=M(\nu,n+2,x).$
  • We write that the determinant of the resulting system is zero (since $(X,Y)\neq(0,0)$, according to Remark 3.7) and we obtain the desired equation.
From the recursion relation [11, (13.3.4)],
$xM(\nu+1,n+3,x)=(n+2)M(\nu+1,n+2,x)-(n+2)M(\nu,n+2,x)$ (3.7)
and from [11, (13.3.3)]
$\nu M(\nu+1,n+2,x)=(n+1)M(\nu,n+1,x)+(\nu-n-1)M(\nu,n+2,x).$ (3.8)
Inserting Equation (3.8) in Equation (3.5), we get
$(n+1)(n+x)X+2x(\nu-n-1)Y=0.$ (3.9)
Inserting Equation (3.7) in Equation (3.6), we get
$02\nu M(\nu+1,n+2,x)+(n+1-2\nu-x)Y=$ 0 .
Inserting Equation (3.8) in the previous equation, we obtain
$02(n+1)X-(n+1+x)Y=$ 0 . (3.10)
The system formed by Equations (3.9) and (3.10) is singular if and only if
$\left|\begin{array}[]{cc}(n+1)(n+x)&2x(\nu-n-1)\\ 2(n+1)&-(n+1+x)\end{array}\right|=0.$
This gives the equation
$x^{2}+(4\nu-2n-3)x+n(n+1)=0,$ (3.11)
which can be rewritten
$x^{2}-(2\eta+2n+1)x+n(n+1)=0.$ (3.12)
We revert to the variable $\beta$ and we obtain
$\beta^{2}-2(2\eta+2n+1)\beta+4n(n+1)=0\,.$ (3.13)
A similar proof of the same formula is given in [23], using Whittaker functions rather than Kummer functions.
Solving for $\beta$ and using again the necessary condition $\beta>2n$, we finally obtain Saint-James Formula [^7] Saint-James [28] gives this formula in a footnote, writing $h$ for $\beta$ and $\lambda$ for $\eta$, with the following justification “En utilisant les propriétés des fonctions hypergéométriques confluentes”. He probably had to prove first (3.13) but this equation does not appear explicitly in [28].:
$\beta=2\eta+2n+1+\sqrt{(2\eta+1)^{2}+8\eta n}\,.$ (3.14)
In addition, since $\eta>0$, Equation (3.14) immediately implies
$\beta>2(n+1).$
This completes the proof of Theorem 1.1.
Remark 3.8.
We can rewrite (3.13) in the form
$(\beta-(2n+1))^{2}=4\lambda+1\,,$
which holds at the crossing point for $G_{n,\beta}$ and $G_{n+1,\beta}$, with $\eta=\beta\lambda$.
We then get an alternative form of Saint-James formula:
$\beta=(2n+1)+\sqrt{4\,\lambda+1}\,.$ (3.15)
3.5 Interlacing between intersections and minima
We study in this section the relation between the points of intersection and the minima $\beta_{\min}(n)$. Our results give in particular a proof of Theorem 1.2. The derivatives are again taken with respect to $\beta$.
Proposition 3.9.
Let $n\in\mathbb{N}$ and let $\beta^{*}$ be a solution in $(0,\infty)$ of the equation
$\eta(n,\beta)=\eta(n+1,\beta).$
Then,
$\eta^{\prime}(n,\beta^{*})>0$
and
$\eta^{\prime}(n+1,\beta^{*})<0.$
Proof.
We use the notation $\eta^{*}_{n}=\eta(n,\beta^{*})=\eta(n+1,\beta^{*})$ and $x^{*}:=\beta^{*}/2$ to make the formulas more readable. A direct computation, using Point (i) of Proposition 3.3, gives
${\eta^{\prime}(n,\beta^{*})}=c_{n}\frac{(x^{*})^{2}-(2\eta^{*}_{n}+2n)x^{*}+n^{2}}{2x^{*}}$
and
${\eta^{\prime}(n+1,\beta^{*})}=c_{n+1}\frac{(x^{*})^{2}-(2\eta^{*}_{n}+2n+2)x^{*}+(n+1)^{2}}{2x^{*}},$
with $c_{n}$ and $c_{n+1}$ positive factors. Taking into account that $n$, $x^{*}$ and $\eta^{*}$ satisfy Equation (3.12), we obtain
${\eta^{\prime}(n,\beta^{*})}=c_{n}\frac{x^{*}-n}{2x^{*}}$
and
${\eta^{\prime}(n+1,\beta^{*})}=c_{n+1}\frac{n+1-x^{*}}{2x^{*}}.$
According to Theorem 1.1, $x^{*}>n+1$, and therefore $\eta^{\prime}(n,\beta^{*})>0$ and $\eta^{\prime}(n+1,\beta^{*})<0$.
The following theorem is now easy to prove, and implies Theorem 1.2. We use the convention $\beta_{\min}(0)=0$.
Theorem 3.10.
For any $n\in\mathbb{N}$, there exists a unique solution $\beta_{n}$ in $[0,+\infty)$ of the equation
$\lambda(n,\beta)=\lambda(n+1,\beta).$
In addition,
  1. (i)
    the sequences $(\beta_{\min}(n))$ and $(\beta_{n})$ are both increasing,
  2. (ii)
    for any $n\in\mathbb{N}$, $\beta_{\min}(n)<\beta_{n}<\beta_{\min}(n+1)$,
  3. (iii)
    $\lambda(\beta)=\lambda(0,\beta)$ when $\beta\in[0,\beta_{0}]$ and, for $n\in\mathbb{N}\setminus\{0\}$, $\lambda(\beta)=\lambda(n,\beta)$ when $\beta\in[\beta_{n-1},\beta_{n}]$. Furthermore, if $\beta\notin\{\beta_{n}\,:\,n\in\mathbb{N}\}$, there is a unique $k\in\mathbb{N}$ such that $\lambda(\beta)=\lambda(k,\beta)$.
Proof.
Let $\beta^{*}$ be a solution of
$\eta(n,\beta)=\eta(n+1,\beta).$
The signs of the derivatives established in Proposition 3.9 imply that
$\beta_{\min}(n)<\beta^{*}<\beta_{\min}(n+1)$
(the case $n=0$ is included using the convention that $\beta_{\min}(0)=0$). Let us note that, since at least one solution exists according to Proposition 3.6, we have shown $\beta_{\min}(n)<\beta_{\min}(n+1)$.
By the same argument, any other solution must also belong to the interval
$(\beta_{\min}(n),\beta_{\min}(n+1)).$
Since $\beta\mapsto\eta(n,\beta)$ is strictly increasing and $\beta\mapsto\eta(n+1,\beta)$ strictly decreasing in the interval $(\beta_{\min}(n),\beta_{\min}(n+1))$, there cannot be another solution. We now use the notation $\beta_{n}:=\beta^{*}$ and we note that we have proved Point (ii). Since the sequence $(\beta_{\min}(n))$ is increasing, so is the sequence $(\beta_{n})$, which proves Point (i).
To prove Point (iii), we remark that multiplying the values $\eta(n,\beta)$ by $\beta$ does not change their ordering, so that $\lambda(n,\beta)<\lambda(n+1,\beta)$ for $\beta<\beta_{n}$ and $\lambda(n,\beta)>\lambda(n+1,\beta)$ for $\beta>\beta_{n}$. Combining this with the fact that the sequence $(\beta_{n})$ is increasing, we first deduce that, for $\beta\in(0,\beta_{0})$,
$\lambda(0,\beta)<\lambda(1,\beta)<\lambda(2,\beta)<\dots\,,$
and in particular $\lambda(\beta)=\lambda(\beta,0)$. Then, for any $n\in\mathbb{N}\setminus\{0\}$ and $\beta\in(\beta_{n-1},\beta_{n})$, we find
$\lambda(n,\beta)<\lambda(n+1,\beta)<\lambda(n+2,\beta)<\dots$
and
$\lambda(0,\beta)>\lambda(1,\beta)>\dots>\lambda(n-1,\beta)>\lambda(n,\beta).$
In particular, $\lambda(\beta)=\lambda(n,\beta)$. This establishes Point (iii) when $\beta$ is not a point of intersection. The full statement follows by continuity.
Combining Point (iii) of the previous theorem with the inequality $\beta_{n-1}>2n$ (deduced from the Saint-James Formula (3.14)) and with Lemma 3.5, we obtain, for all $\beta>0$,
$\lambda(\beta)<\beta\,.$
This inequality is proved for a general domain in $\mathbb{R}^{2}$ by Colbois-Léna-Provenzano-Savo [7].
Let us now recall the definition
$\eta_{n}^{*}:=\eta(n,\beta_{n})=\eta(n+1,\beta_{n})$
appearing in the statement of Conjecture 1.5.
Since $\eta(\beta)=\eta(n+1,\beta)$ for $\beta\in[\beta_{n},\beta_{n+1}]$ and since $\beta\mapsto\eta(n+1,\beta)$ is decreasing in $[\beta_{n},\beta_{\min}(n+1)]$ and increasing in $[\beta_{\min}(n+1),\beta_{n+1}]$, we have
$\eta(n+1,\beta)\leq\max\{\eta_{n}^{*},\eta_{n+1}^{*}\}\mbox{ for all }\beta\in[\beta_{n},\beta_{n+1}]\,.$ (3.16)
In addition,
$\eta(\beta)\leq\eta_{0}^{*}\mbox{ for all }\beta\in(0,\beta^{*}_{0}]\,,$ (3.17)
since $\eta(\beta)=\eta(0,\beta)$ and $\beta\mapsto\eta(0,\beta)$ is strictly increasing in $(0,\beta_{0}^{*}]$.
Conjecture 1.3 is therefore equivalent to
$\eta^{*}_{n}<\Theta_{0}\mbox{ for all }n\in\mathbb{N}.$ (3.18)
From known asymptotic results (1.5), $\eta_{n}^{*}$ tends to $\Theta_{0}$. Therefore, it would be enough to prove that the sequence $(\eta^{*}_{n})$ is strictly increasing to obtain Inequality (3.18). This shows that Conjecture 1.4 implies Conjecture 1.3.
Notice also that we can deduce from (3.13) that if Conjecture 1.5 holds, i. e. if $\lambda(\beta)$ is increasing then
$\beta_{n+1}\geq 2+\beta_{n}\,.$ (3.19)
In particular, this holds for $n$ large enough, using Fournais-Helffer [13] monotonicity result, which will be recalled in Section 4 and analyzed in more detail. This will also be confirmed numerically, for $n\leq 400$.
4 Asymptotic results for the disk revisited
Starting from the asymptotic analysis given in [14] we show that these asymptotics could be refined at the intersection points i.e. for the sequence $\beta_{n}$ as $n\rightarrow+\infty$. We then show how the comparison with the information coming from Saint-James formula leads to complete asymptotics of $\beta_{n}$ and $\lambda(\beta_{n})=\lambda(n,\beta_{n})$. This validates asymptotically a conjecture on the monotonicity of $\beta_{n+1}-\beta_{n}$.
4.1 The case of the disk (reminder after [13])
With (1.3) in mind, we also recall that $\xi_{0}<0$ is the point where $\xi\mapsto\lambda^{DG}(\xi)$ attains its unique (non degenerate) minimum and that $C_{1}$ is given by
$C_{1}=\frac{u_{0}^{2}(0)}{3}\;.$ (4.1)
where $u_{0}=\phi_{\xi_{0}}$, is the positive, normalized eigenfunction associated with $\lambda^{DG}(\xi_{0})$.
Numerically (see [5], [13] or [27])
$C_{1}\sim 0.254\,,\,\xi_{0}\sim-0.768\,.$
We also recall that
$\Theta_{0}=\xi_{0}^{2}\,.$ (4.2)
Theorem 4.1 (Eigenvalue asymptotics for the disk).
Suppose that $\mathbb{D}$ is the unit disk. Define $\delta(m,\beta)$, for $m\in{\mathbb{Z}}$, $\beta>0$, by
$\delta(m,\beta)=m-\tfrac{\beta}{2}-\xi_{0}\sqrt{\beta}.$ (4.3)
Then there exist (computable) constants $C_{0},\delta_{0}\in{\mathbb{R}}$ such that if
$\Delta_{\beta}=\inf_{m\in{\mathbb{Z}}}|\delta(m,\beta)-\delta_{0}|\;,$ (4.4)
then
$\lambda(\beta)=\Theta_{0}\beta-C_{1}\sqrt{\beta}+3C_{1}\sqrt{\Theta_{0}}\big{(}\Delta_{\beta}^{2}+C_{0}\big{)}+{\mathcal{O}}(\beta^{-\frac{1}{2}})\;.$ (4.5)
The asymptotics (4.5) is deduced (see the proof in [14] or a more detailed analysis in the next subsection) from the asymptotics
$\lambda(m,\beta)=\Theta_{0}\beta-C_{1}\sqrt{\beta}+3C_{1}\sqrt{\Theta_{0}}\big{(}|\delta(m,\beta)-\delta_{0}|^{2}+C_{0}\big{)}+{\mathcal{O}}(\beta^{-\frac{1}{2}})\,,$ (4.6)
and then considering the infimum.
Although not used here, notice that the $3D$-case was analyzed by Helffer-Morame [21] (see also [15, Chap. 9]) and, for the version with three terms, in [17].
4.2 Theorem 4.1 revisited
We now revisit the proof of Theorem 4.1 as given in [14] (see also [17]) in order to improve its conclusions when considering $\beta=\beta_{n}$.
We start with the Neumann operator
$\mathfrak{h}_{0}(\zeta)=-\frac{d^{2}}{d\tau^{2}}+(\tau+\zeta)^{2}$
on $L^{2}({\mathbb{R}}_{+},d\tau)\;.$
Let $D(t)=\{x\in{\mathbb{R}}^{2}\,|\,|x|\leq t\}$ be the disc with radius $t$. Let $\widetilde{Q}_{\beta}$ be the quadratic form
[…]
with domain $\{u\in H^{1}(D(1)\setminus D(\frac{1}{2}))\,|\,u(x)=0\text{ on }|x|=\frac{1}{2}\}$. Let $\tilde{\lambda}(\beta)$ be the lowest eigenvalue of the corresponding self-adjoint operator (Friedrichs extension). Using the Agmon estimates in the normal direction, it can be proved
$\lambda(\beta)=\tilde{\lambda}(\beta)+{\mathcal{O}}(\beta^{-\infty})\;.$ (4.7)
By changing to boundary coordinates (if $(r,\theta)$ are usual polar coordinates, then $t=1-r$, $s=\theta$), the quadratic form $\widetilde{Q}_{\beta}[u]$ becomes,
[…]
Performing the scaling $\tau=\sqrt{\beta}\,t$ and decomposing in Fourier modes,
$u=\sum_{m}e^{ims}\phi_{m}(t)\;,$
leads to
$\tilde{\lambda}(\beta)=\beta\inf_{m\in{\mathbb{Z}}}e_{\delta(m,\beta),\beta}\;.$ (4.9)
Here the function $\delta(m,\beta)$ was defined in (4.3) and $e_{\delta,\beta}$ is the lowest eigenvalue of the operator $\mathfrak{h}(\delta,\beta)$ associated with the quadratic form $q_{\delta,\beta}$ on $L^{2}((0,\sqrt{\beta}/2);(1-\beta^{-1/2}\tau))d\tau)$ (with Neumann boundary condition at 00 and Dirichlet at $\sqrt{\beta}/2$):
$q_{\delta,\beta}[\phi]=\int_{0}^{\sqrt{\beta}/2}$ $(1-\tfrac{\tau}{\sqrt{\beta}})^{-1}\big{(}(\tau+\xi_{0})+\beta^{-\frac{1}{2}}(\delta-\tfrac{\tau^{2}}{2})\big{)}^{2}$
$+(1-\tfrac{\tau}{\sqrt{\beta}})|\phi^{\prime}(\tau)|^{2}\,d\tau\;.$ (4.10)
The self-adjoint Neumann operator ${\mathfrak{h}}(\delta,\beta)$ associated with $q_{\delta,\beta}$ (on the space $L^{2}((0,\sqrt{\beta}/2);(1-\beta^{-1/2}\tau)d\tau)$) is
${\mathfrak{h}}(\delta,\beta)$ $=-(1-\tfrac{\tau}{\sqrt{\beta}})^{-1}\frac{d}{d\tau}(1-\tfrac{\tau}{\sqrt{\beta}})\frac{d}{d\tau}$
$\quad+(1-\tfrac{\tau}{\sqrt{\beta}})^{-2}\big{(}(\tau+\xi_{0})+\beta^{-\frac{1}{2}}(\delta-\tfrac{\tau^{2}}{2})\big{)}^{2}\;.$ (4.11)
We will only consider $\delta$ varying in a fixed bounded set. In this case, we know that there exists a $d>0$ such that if $\beta>d^{-1}$, then the spectrum of $\mathfrak{h}(\delta,\beta)$ contained in $(-\infty,\Theta_{0}+d)$ consists of exactly one simple eigenvalue.
We can formally develop ${\mathfrak{h}}(\delta,\beta)$ as
${\mathfrak{h}}(\delta,\beta)={\mathfrak{h}}_{0}+\beta^{-\frac{1}{2}}{\mathfrak{h}}_{1}+\beta^{-1}{\mathfrak{h}}_{2}+{\mathcal{O}}(\beta^{-\frac{3}{2}})\;.$
with
${\mathfrak{h}}_{0}$ $=-\frac{d^{2}}{d\tau^{2}}+(\tau+\xi_{0})^{2}\;,$
${\mathfrak{h}}_{1}$ $=\frac{d}{d\tau}+2(\tau+\xi_{0})(\delta-\tfrac{\tau^{2}}{2})+2\tau(\tau+\xi_{0})^{2}\;,$
${\mathfrak{h}}_{2}$ $=\tau\frac{d}{d\tau}+(\delta-\tfrac{\tau^{2}}{2})^{2}+4\tau(\tau+\xi_{0})(\delta-\tfrac{\tau^{2}}{2})+3\tau^{2}(\tau+\xi_{0})^{2}\;.$ (4.12)
Let $u_{0}$ be the ground state eigenfunction of ${\mathfrak{h}}_{0}$ with eigenvalue $\Theta_{0}$, where ${\mathfrak{h}}_{0}$ is considered as a selfadjoint operator on $L^{2}({\mathbb{R}}_{+};d\tau)$ with Neumann boundary condition at 00. Let $R_{0}$ be the regularized resolvent, which is defined by
$R_{0}\phi=\begin{cases}({\mathfrak{h}}_{0}-\Theta_{0})^{-1}\phi\;,&\phi\perp u_{0}\;,\\ \quad 0\;,&\phi\parallel u_{0}\,.\end{cases}$
Here the orthogonality is measured with respect to the usual inner product in $L^{2}({\mathbb{R}}_{+};d\tau)$.
Let $\lambda_{1}$ and $\lambda_{2}$ be given by
$\lambda_{1}$ $:=\langle u_{0}\,|\,{\mathfrak{h}}_{1}u_{0}\rangle\;,$ $\lambda_{2}$ $:=\lambda_{2,1}+\lambda_{2,2}\;,$
$\lambda_{2,1}$ $:=\langle u_{0}\,|\,{\mathfrak{h}}_{2}u_{0}\rangle\;,$ $\lambda_{2,2}$ $:=\langle u_{0}\,|\,({\mathfrak{h}}_{1}-\lambda_{1})u_{1}\rangle\;,$ (4.13)
Here the inner products are the usual inner products in $L^{2}({\mathbb{R}}_{+};d\tau)$. The functions $u_{1},u_{2}$ are given as
$u_{1}$ $=-R_{0}({\mathfrak{h}}_{1}-\lambda_{1})u_{0}\;,$ $u_{2}$ $=-R_{0}\big{\{}({\mathfrak{h}}_{1}-\lambda_{1})u_{1}+({\mathfrak{h}}_{2}-\lambda_{2})u_{0}\big{\}}\;.$ (4.14)
Notice that $u_{0}\in{\mathcal{S}}(\overline{{\mathbb{R}}_{+}})$ and that $R_{0}$ maps ${\mathcal{S}}(\overline{{\mathbb{R}}_{+}})$ (continuously) into itself. Therefore, $u_{0},u_{1},u_{2}$ (and their derivatives) are rapidly decreasing functions on ${\mathbb{R}}_{+}$.
Let $\chi\in C_{0}^{\infty}({\mathbb{R}})$ be a usual cut-off function, such that
$\chi(t)$ $=1\quad\text{ for }|t|\leq\tfrac{1}{8}\;,$ ${\rm supp}\chi$ $\subset[-\tfrac{1}{4},\tfrac{1}{4}]\;,$ (4.15)
and let $\chi_{\beta}(\tau)=\chi(\tau\beta^{-\frac{1}{4}})$
Our trial state is defined by
$\psi:=\chi_{\beta}\big{\{}u_{0}+\beta^{-\frac{1}{2}}u_{1}+\beta^{-1}u_{2}\big{\}}\;.$ (4.16)
A calculation (using in particular the exponential decay of the involved functions) gives that
$\big{\|}\big{\{}{\mathfrak{h}}(\delta,\beta)-\big{(}\Theta_{0}+\lambda_{1}\beta^{-\frac{1}{2}}+\lambda_{2}\beta^{-1}\big{)}\big{\}}\psi\big{\|}_{L^{2}([0,\sqrt{\beta}/2];(1-\beta^{-1/2}\tau)d\tau)}={\mathcal{O}}(\beta^{-\frac{3}{2}})\;,$
$\|\psi\|_{L^{2}([0,\sqrt{\beta}/2];(1-\beta^{-1/2}\tau)d\tau)}=1+{\mathcal{O}}(\beta^{-\frac{1}{2}})\;,$ (4.17)
where the constant in ${\mathcal{O}}$ is uniform for $\delta$ in bounded sets.
Therefore, we have proved that (uniformly for $\delta$ varying in bounded sets)
$e_{\delta,\beta}=\Theta_{0}+\lambda_{1}\beta^{-\frac{1}{2}}+\lambda_{2}\beta^{-1}+{\mathcal{O}}(\beta^{-\frac{3}{2}})\;.$ (4.18)
It remains to calculate $\lambda_{1},\lambda_{2}$ and, in particular, deduce their dependence on $\delta$.
A standard calculation (which can for instance be found in [14, Section 2]) gives that
$\lambda_{1}=-C_{1}\;.$ (4.19)
It was harder to calculate $\lambda_{2}$ explicitly but it is proven in [14] that
$\lambda_{2}(\delta)=3C_{1}\sqrt{\Theta_{0}}\big{(}(\delta-\delta_{0})^{2}+C_{0}\big{)}\;.$ (4.20)
Remembering (4.7) and (4.9) this finishes the proof of Theorem 4.1. A computation of $\delta_{0}$ will be given later.
The same proof (but pushing the expansion further as done for instance in [17]) gives the following extension
Proposition 4.2.
For any $2\leq N\in\mathbb{N}$, uniformly for $\delta$ varying in bounded sets, we have
$e_{\delta,\beta}=\Theta_{0}+\lambda_{1}\beta^{-\frac{1}{2}}+\lambda_{2}(\delta)\beta^{-1}+\sum_{j=3}^{N}\lambda_{j}(\delta)\beta^{-j/2}+{\mathcal{O}}(\beta^{-\frac{N+1}{2}})\;.$ (4.21)
Moreover the $\lambda_{j}(\delta)$ are polynomials of $\delta$.
Proof.
Following the previous proof (corresponding to $N=2$)
$\psi_{N}:=\chi_{\beta}\big{\{}\sum_{j=0}^{N}\beta^{-\frac{j}{2}}u_{j}\}$ (4.22)
and
$\big{\|}\big{\{}{\mathfrak{h}}(\delta,\beta)-\big{(}\Theta_{0}+\lambda_{1}\beta^{-\frac{1}{2}}+\sum_{j=2}^{N}\lambda_{j}(\delta)\beta^{-j/2}\big{)}\big{\}}\psi_{N}\big{\|}_{L^{2}([0,\sqrt{\beta}/2];(1-\beta^{-1/2}\tau)d\tau)}={\mathcal{O}}(\beta^{-\frac{N+1}{2}})\;,$
$\|\psi_{N}\|_{L^{2}([0,\sqrt{\beta}/2];(1-\beta^{-1/2}\tau)d\tau)}=1+{\mathcal{O}}(\beta^{-\frac{1}{2}})\;,$ (4.23)
where the constant in ${\mathcal{O}}$ is uniform for $\delta$ in bounded sets.
Note for later that the normalized eigenfunction $u_{\delta,\beta}$ associated with $e_{\delta,\beta}$ satisfies, for any interval $(0,M)$ and any $N$,
$||u_{\delta,\beta}-u_{0}-\sum_{j=1}^{N}\beta^{-\frac{j}{2}}\hat{u}_{j}||_{H^{1}((0,M);(1-\beta^{-1/2}\tau)d\tau)}\leq C_{M,N}\beta^{-(N+1)/2}\,,$ (4.24)
for $\beta$ large enough.
Here the $\hat{u}_{j}$ are in $\mathcal{S}(\overline{\mathbb{R}_{+}})$ and depend polynomially of $\delta$.
Combining with (4.7), (4.9) and considering for $n$ large $\beta=\beta_{n}$, we obtain
Corollary 4.3.
For any $N$, we have with $\beta=\beta_{n}$ and $\delta=\delta(n,\beta_{n})$
${\lambda(\beta_{n})}=\lambda(n,\beta_{n})=\beta\Big{(}\Theta_{0}+\lambda_{1}\beta^{-\frac{1}{2}}+\lambda_{2}(\delta)\beta^{-1}+\sum_{j=3}^{N}\lambda_{j}(\delta)\beta^{-j/2}+{\mathcal{O}}(\beta^{-\frac{N+1}{2}})\Big{)}\,.$ (4.25)
Theorem 4.4.
There exist sequences $\hat{\delta}_{j}$ ($j\geq 1$) and $\hat{\lambda}_{j}$ ($j\geq 3$) such that, for any $N$, we have
$\delta(n,\beta_{n})=\delta_{0}-\frac{1}{2}+\sum_{1\leq j\leq N}\hat{\delta}_{j}\,\beta_{n}^{-j/2}+{\mathcal{O}}(\beta_{n}^{-\frac{N+1}{2}})\,,$ (4.26)
and
$\lambda_{1}(\beta_{n})/\beta_{n}=\Theta_{0}+\lambda_{1}\,\beta_{n}^{-\frac{1}{2}}+\lambda_{2}(\delta_{0}-\frac{1}{2})\,\beta_{n}^{-1}+\sum_{j=3}^{N}\check{\lambda}_{j}\,\beta_{n}^{-j/2}+{\mathcal{O}}(\beta_{n}^{-\frac{N+1}{2}})\,.$ (4.27)
To get (4.26), we simply compare the expansions given for $\lambda_{1}(n,\beta_{n})$ (i.e $\delta=\delta(n,\beta_{n})$ and $\lambda_{1}(n+1,\beta_{n})$ (i.e. $\delta=\delta(n+1,\beta_{n})$).
Let us describe the first step in more detail. We use (4.5) for $\beta=\beta_{n}$:
$\lambda(\beta_{n})/\beta_{n}=\Theta_{0}-C_{1}(\beta_{n})^{-1/2}+3C_{1}\Theta_{0}^{1/2}\big{(}(\delta(n,\beta_{n})-\delta_{0})^{2}+C_{0}\big{)}(\beta_{n})^{-1}+\mathcal{O}(\beta_{n}^{-3/2})\,.$ (4.28)
At the crossing we get, by using (4.6) for $\beta=\beta_{n}$ and for both $m=n$ and $m=n+1$,
$\delta(n,\beta_{n})+\delta(n+1,\beta_{n})=2\delta_{0}+\mathcal{O}(\beta_{n}^{-1/2})\,.$ (4.29)
which implies using (4.3), (4.26) and
$\delta_{0}-\delta(n,\beta_{n})=\frac{1}{2}+\mathcal{O}(\beta_{n}^{-1/2})\,.$ (4.30)
Hence we obtain at the crossing point $\beta_{n}$
$\frac{\lambda(\beta_{n})}{\beta_{n}}:=\eta_{n}^{*}=\Theta_{0}-C_{1}(\beta_{n})^{-1/2}+3C_{1}\Theta_{0}^{1/2}(\frac{1}{4}+C_{0})(\beta_{n})^{-1}+\mathcal{O}(\beta_{n}^{-3/2})\,.$ (4.31)
All these formulas correspond to an asymptotic as $n\rightarrow+\infty$.
By recursion we get expansions at any order.
Remark 4.5.
Combining with (3.15) at $\beta_{n}$ we get, using (4.26) and (4.27),
$\delta_{0}=\frac{1}{2}\Theta_{0}^{-1/2}C_{1}\sim 0.0975\,.$ (4.32)
This formula seems to be new and was not obtained in [14].
Notice that we have not at the moment mixed these results with the Saint-James formula.
4.3 Implementation of Saint-James formula
In this subsection, we give a complete expansion for $\beta_{n}$ as an asymptotic series in powers of $n^{-1/2}$.
Theorem 4.6.
There is an infinite sequence $\hat{\kappa}_{j}$ ($j\geq 0$) such that
$\beta_{n}=2n+\xi_{1}n^{\frac{1}{2}}+\hat{\kappa}_{0}+\sum_{j\geq 1}\hat{\kappa}_{j}n^{-j/2}+\mathcal{O}(n^{-\infty})\,,$ (4.33a)
with
$\xi_{1}=-2^{3/2}\xi_{0}\,,\,\hat{\kappa}_{0}=1-2\delta_{0}+2\xi_{0}^{2}\,.$ (4.33b)
In particular, this implies
Corollary 4.7.
$\gamma_{n}:=\beta_{n+1}-\beta_{n}=2+\frac{\xi_{1}}{2}n^{-1/2}+\hat{\gamma}_{3}n^{-3/2}+\mathcal{O}(n^{-2})\,.$ (4.34)
and
$\gamma_{n+1}-\gamma_{n}=-\frac{1}{4}\xi_{1}n^{-3/2}+\mathcal{O}(n^{-5/2})\,.$ (4.35)
This in particular proves for $n$ large that the sequence $\gamma_{n}$ is decreasing.
Proof of Theorem 4.6
To prove (4.33), we implement our candidate for the expansion into (3.15) which reads at $\beta_{n}$:
$\beta_{n}-(2n+1)=\sqrt{4\lambda(n,\beta_{n})+1}\,.$
The $\hat{\kappa}_{j}$ are then obtained by recursion.
Let us more explicitly describe the first step.
We start from (3.15) at $\beta=\beta_{n}$ but we implement (4.27)
$\lambda(\beta_{n})=\beta_{n}\big{(}\Theta_{0}-C_{1}(\beta_{n})^{-1/2}+3C_{1}\Theta_{0}^{1/2}(\frac{1}{4}+C_{0})(\beta_{n})^{-1}\big{)}+\mathcal{O}(n^{-1/2})\,.$ (4.36)
This leads to
$4\lambda(\beta_{n})+1=4\beta_{n}\big{(}\Theta_{0}-C_{1}(\beta_{n})^{-1/2}+C_{2}(\beta_{n})^{-1}\big{)}+\mathcal{O}(n^{-1/2})\,,$
with
$C_{2}:=3C_{1}\Theta_{0}^{1/2}(\frac{1}{4}+C_{0})+\frac{1}{4}\,.$
We then get
$\sqrt{4\lambda(\beta_{n})+1}=2\Theta_{0}^{\frac{1}{2}}\beta_{n}^{1/2}\big{(}1-C_{1}\Theta_{0}^{-1}\beta_{n}^{-1/2}+C_{2}\Theta_{0}^{-1}\beta_{n}^{-1}+\mathcal{O}(n^{-3/2})\big{)}^{\frac{1}{2}}$
which implies
$\sqrt{4\lambda(\beta_{n})+1}=2\Theta_{0}^{\frac{1}{2}}\beta_{n}^{1/2}-C_{1}\Theta_{0}^{-1/2}+C_{3}\Theta_{0}^{-1}\beta_{n}^{-1/2}+\mathcal{O}(n^{-1})\,,$
with $C_{3}$ computable.
This leads to
$\beta_{n}=2n+1+2\Theta_{0}^{\frac{1}{2}}\beta_{n}^{1/2}-C_{1}\Theta_{0}^{-1/2}+C_{3}\Theta_{0}^{-1}\beta_{n}^{-1/2}+\mathcal{O}(n^{-1})\,.$
Implementing $\beta_{n}=2n+\xi_{1}n^{\frac{1}{2}}+\hat{\kappa}_{0}+\hat{\kappa}_{1}n^{-1/2}+\mathcal{O}(n^{-1})$, we get the existence of explicit $\hat{\kappa}_{0}$ (as given in the statement) and $\hat{\kappa}_{1}$.
5 Strong diamagnetism revisited
5.1 Reminder and first approach
In the case of the disk, it was also proved in [13] the following statement called “strong diamagnetism”.
Theorem 5.1.
Let $\mathbb{D}$ be the disk. Then the left- and right-hand derivatives $\lambda_{\pm}^{\prime}(\beta)$ exist and satisfy
$\lambda_{+}^{\prime}(\beta)\leq\lambda_{-}^{\prime}(\beta)\;,$
$\liminf_{\beta\rightarrow+\infty}\lambda_{+}^{\prime}(\beta)\geq\Theta_{0}-\tfrac{3}{2}C_{1}|\xi_{0}|>0\;.$ (5.1)
In particular, $\beta\mapsto\lambda(\beta)$ is strictly increasing for large $\beta$.
Notice that the strong diamagnetism in dimension 3 is considered by Soeren Fournais and Mikael Persson Sundqvist in [17]. Our analysis of the Saint-James picture has given a more precise description by determining the sequence $(\beta_{n})$ where the left and right derivatives could differ. It is now interesting to combine the informations given by the two approachs.
But let us come back to $\lambda(n,\beta)$ since we are interested in the asymptotics of $\lambda^{\prime}(n,\beta)$. Hence we start from
$\lambda(n,\beta)=\Theta_{0}\beta+\lambda_{1}\beta^{\frac{1}{2}}+\lambda_{2}(\delta(n,\beta)))+\sum_{j=3}^{N}\lambda_{j}(\delta(n,\beta))\beta^{1-j/2}+{\mathcal{O}}(\beta^{1-\frac{N+1}{2}})\,,$
which is a consequence of (4.7), (4.9) and (4.21).
We rewrite this as
$\lambda(n,\beta)=g_{N}(\beta)+r_{N}(\beta)\,,$
with
$g_{N}(\beta):=\Theta_{0}\beta+\lambda_{1}\beta^{\frac{1}{2}}+\lambda_{2}(\delta(n,\beta)))+\sum_{j=3}^{N}\lambda_{j}(\delta(n,\beta))\beta^{1-j/2}\,.$
The problem is that it could be difficult to prove directly that we can differentiate the expansion, since we have no good control of $r^{\prime}_{N}(\beta)$.
One way to avoid this difficulty is to use the observation in [13] that, for any $\epsilon>0$,
$\lambda_{+}^{\prime}(\beta)$ $=-2\Re\langle\phi_{{+}}(\beta)\,|\,{\bf A}\cdot(-i\nabla-\beta{\bf A})\phi_{{+}}(\beta)\rangle$
$=\frac{1}{\epsilon}\langle\phi_{{+}}(\beta)\,|\,\big{(}H_{\beta+\epsilon}-H_{\beta}-\epsilon^{2}{\bf A}^{2}\big{)}\phi_{{+}}(\beta)\rangle\;,$
where $\phi_{+}(\gamma)$ is a suitable eigenfunction depending analytically of $\gamma$ around $\beta$ and associated with $\lambda(\gamma)$ for $\gamma\in[\beta,\beta+\epsilon)$. Therefore, the variational principle implies
$\lambda_{+}^{\prime}(\beta)\geq\frac{\lambda(\beta+\epsilon)-\lambda(\beta)}{\epsilon}-\epsilon\|{\bf A}\|_{L^{\infty}(\Omega)}^{2}\,,$
and we have trivially
$\lambda_{+}^{\prime}(\beta)\leq\frac{\lambda(\beta+\epsilon)-\lambda(\beta)}{\epsilon}\,.$
For $\epsilon>0$ small enough, it is consequently sufficient to have an expansion of $\frac{\lambda(\beta+\epsilon)-\lambda(\beta)}{\epsilon}$.
For $\beta=\beta_{n}$, we have
$\frac{\lambda(\beta_{n}+\epsilon)-\lambda(\beta_{n})}{\epsilon}=\epsilon^{-1}\Big{(}g_{N}(\beta_{n}+\epsilon))-g_{N}(\beta_{n})\Big{)}+\epsilon^{-1}{\mathcal{O}}((\beta_{n})^{1-\frac{N+1}{2}})$
But we can control $g_{N}(\beta)$ and its derivatives (note that $\delta^{\prime}(m,\beta)=\mathcal{O}(1)$, $\delta^{\prime\prime}=\mathcal{O}(\beta^{-3/2})$,…..)
Using its Taylor expansion, with remainder term and taking $N$ large and $\epsilon=(\beta_{n})^{-N/2}$, we can justify the differentiation term by term of the expansion of $\lambda(n,\beta)$. In particular we get:
$\lambda_{+}^{\prime}(\beta_{n})=\Theta_{0}+\frac{3}{2}C_{1}|\xi_{0}|+\check{C}_{2}(\beta_{n})^{-1/2}+\mathcal{O}((\beta_{n})^{-1})\,.$ (5.2)
End of the proof.
We only give the details for the first step. We recall that
$\delta_{0}-\delta(n,\beta_{n})=\frac{1}{2}+\delta_{1}n^{-1/2}+\mathcal{O}(n^{-1})\,.$ (5.3)
Now we can differentiate (4.6). So we get
$\lambda^{\prime}(m,\beta)=\Theta_{0}-C_{1}\beta^{-1/2}+6C_{1}\sqrt{\Theta_{0}}\big{(}\delta(m,\beta)-\delta_{0})\delta^{\prime}(m,\beta)\big{)}+\mathcal{O}(\beta^{-1/2})\,,$ (5.4)
Observing that
$\delta^{\prime}(m,\beta)=-\frac{1}{2}-\frac{1}{2}\xi_{0}\beta^{-1/2}$
we get
$\lambda^{\prime}(m,\beta)=\Theta_{0}-C_{1}\beta^{-1/2}-3C_{1}\sqrt{\Theta_{0}}(\delta(m,\beta)-\delta_{0})(1+\xi_{0}\beta^{-1/2})+\mathcal{O}(\beta^{-1/2})\,,$ (5.5)
We apply this formula with $m=n$, $\beta=\beta_{n}$,
$\lambda^{\prime}(n,\beta_{n})=\Theta_{0}-C_{1}\beta^{-1/2}+3C_{1}\sqrt{\Theta_{0}}(\frac{1}{2}+\hat{\delta}_{1}n^{-1/2}+\mathcal{O}(n^{-1}))(1+\xi_{0}\beta^{-1/2})+\mathcal{O}(\beta^{-1/2})\,,$ (5.6)
$\lambda^{\prime}(n,\beta_{n})=\Theta_{0}+\frac{3}{2}C_{1}\sqrt{\Theta_{0}}+\hat{C}_{1}n^{-1/2}+\mathcal{O}(n^{-1})\,.$ (5.7)
Hence we have a two terms expansion which improves (5.16).
Using the more complete expansion, we get
Theorem 5.2.
We have
$\lambda^{\prime}(n,\beta_{n})=\Theta_{0}+\frac{3}{2}C_{1}\sqrt{\Theta_{0}}+\sum_{j\geq 1}\hat{C}_{1}n^{-j/2}+\mathcal{O}(n^{-\infty})\,.$ (5.8)
We will give an alternative proof in Subsection 5.4.
5.2 About $f_{n,\beta_{n}}$.
We would like to consider the proof of
$\lim_{n\rightarrow+\infty}(\beta_{n})^{-1/2}|f_{n,\beta_{n}}(1)|^{2}=3C_{1}=u_{0}(0)^{2}\,,$
and prove more generally
$(\beta_{n})^{-1/2}|f_{n,\beta_{n}}(1)|^{2}=u(0)^{2}+\sum_{j\geq 1}\check{C}_{j}n^{-j/2}+\mathcal{O}(n^{-\infty})\,.$ (5.9)
This is actually a consequence of the comparison between $f_{n,\beta}$ and the quasi-mode constructed in (4.24). Taking the trace at 00, we get
$u_{\delta(n,\beta),\beta}(0)=u_{0}(0)+\sum_{j=1}^{N}\hat{c}_{j}(\delta)\beta^{-j/2}+\mathcal{O}(\beta^{-(N+1)/2})\,.$
We finally observe that
$f_{n,\beta}(1)=\beta^{1/4}u_{\delta(n,\beta),\beta}(0)+\mathcal{O}(\beta^{-\infty})\,.$
The power of $\beta$ appears when doing the change of variable $t=\beta^{-1/2}\tau$. The error comes from the comparison between the disk and the annulus. We can take $\beta=\beta_{n}$ and the comparison between $\beta_{n}$ and $n$ to obtain a two terms asymptotics.
Using the complete expansion of $\delta(n,\beta_{n})$ we have
Proposition 5.3.
For any $N>0$,
$\beta_{n}^{-1/4}f_{n,\beta_{n}}(1)=u_{0}(0)+\sum_{j\geq 1}^{N}\check{\kappa}_{j}\beta_{n}^{-j/2}+\mathcal{O}(\beta^{-(N+1)/2})\,.$ (5.10)
or
$\beta_{n}^{-1/4}f_{n,\beta_{n}}(1)=u_{0}(0)+\sum_{j\geq 1}^{N}\kappa_{j}n^{-j/2}+\mathcal{O}(n^{-(N+1)/2})\,.$ (5.11)
This in particular proves (5.9).
5.3 Feynman-Hellmann formula in the asymptotic limit
In order to support Conjecture 1.5, let us present some additional results on the variation of $\beta\mapsto\lambda(n,\beta)$. Some of these computations appear already in [19].
Proposition 5.4.
For $n\in\mathbb{N}$, for all $\beta>0$,
$\lambda^{\prime}(n,\beta)=\frac{\lambda(n,\beta)}{\beta}-\frac{1}{2\beta}\left(\lambda(n,\beta)-\big{(}n-\frac{\beta}{2}\big{)}^{2}\right)|f_{n,\beta}(1)|^{2}\,.$ (5.12)
Proof.
Using the notation of Sections 2 and 3, we have, for $\beta>0$,
$\lambda(n,\beta)=\beta\,\eta(n,\beta)$
and
$\lambda^{\prime}(n,\beta)=\eta(n,\beta)+\beta\,\eta^{\prime}(n,\beta)\,.$
Using Point (i) of Proposition 3.3, we get
$\lambda^{\prime}(n,\beta)=\eta(n,\beta)+\frac{\beta}{2}|g_{n,\beta}(\beta^{1/2})|^{2}\left(\left(\frac{n}{\beta^{1/2}}-\frac{\beta^{1/2}}{2}\right)^{2}-\eta(n,\beta)\right).$
Since $g_{n,\beta}(\beta^{1/2})=\beta^{-1/2}f_{n,\beta}(1)$, we obtain the desired formula.
5.4 Application to the strong diamagnetism
We can use Formula (5.12) to evaluate numerically the right derivative of $\beta\mapsto\lambda(\beta)$ at the points of intersection $\beta_{n}$. Indeed, it follows from Theorem 1.2 that
$\lambda^{\prime}_{+}(\beta_{n})=\lambda^{\prime}(n+1,\beta_{n})\,,$ (5.13)
and
$\lambda^{\prime}_{-}(\beta_{n})=\lambda^{\prime}(n,\beta_{n})\,.$ (5.14)
Assuming that Conjecture 1.5 is true, we should find $\lambda^{\prime}_{+}(\beta_{n})\geq 0$.
On the asymptotic side, using the proof in [13] leading to Theorem 5.1, we get
$\lim_{n\rightarrow+\infty}\lambda^{\prime}(n+1,\beta_{n})=\Theta_{0}-\frac{3}{2}C_{1}|\xi_{0}|\,,$ (5.15)
and
$\lim_{n\rightarrow+\infty}\lambda^{\prime}(n,\beta_{n})=\Theta_{0}+\frac{3}{2}C_{1}|\xi_{0}|\,.$ (5.16)
Numerically,
$\Theta_{0}-\frac{3}{2}C_{1}|\xi_{0}|$ $\sim 0.297350,$
$\Theta_{0}+\frac{3}{2}C_{1}|\xi_{0}|$ $\sim 0.882863.$
Let us compute the term which appears in (5.12),
$\frac{1}{2\beta_{n}}\left(\lambda(n,\beta_{n})-\big{(}n-\frac{\beta_{n}}{2}\big{)}^{2}\right)\,.$
We have, using the improvement of (4.26),
$n-\frac{\beta_{n}}{2}=\xi_{0}(\beta_{n})^{1/2}+\delta_{0}-\frac{1}{2}+\mathcal{O}(n^{-1/2})\,,$
and the expansion of $\lambda(n,\beta_{n})$ given in (4.27), we obtain
$\lambda(n,\beta_{n})-\big{(}n-\frac{\beta_{n}}{2}\big{)}^{2}=(-C_{1}-2\xi_{0}(\delta_{0}-\frac{1}{2}))(\beta_{n})^{\frac{1}{2}}+\mathcal{O}(1)\,,$
and
$(\beta_{n})^{-1}\Big{(}\lambda(n,\beta_{n})-\big{(}n-\frac{\beta_{n}}{2}\big{)}^{2}\Big{)}=\xi_{0}(\beta_{n})^{-\frac{1}{2}}+\mathcal{O}(n^{-1})\,.$
Using the asymptotics of $\beta_{n}$, $\lambda(\beta_{n})$ and a complete expansion for $f_{n,\beta_{n}}$, we have a second proof of Theorem 5.2. Using (5.14) and (5.13) we obtain
Theorem 5.5.
We have
$\lambda^{\prime}_{+}(\beta_{n})=\Theta_{0}+\frac{3}{2}C_{1}\sqrt{\Theta_{0}}+\sum_{j\geq 1}\hat{C}_{j}n^{-j/2}+\mathcal{O}(n^{-\infty})\,,$ (5.17)
and
$\lambda^{\prime}_{-}(\beta_{n})=\Theta_{0}-\frac{3}{2}C_{1}\sqrt{\Theta_{0}}+\sum_{j\geq 1}\tilde{C}_{j}n^{-j/2}+\mathcal{O}(n^{-\infty})\,.$ (5.18)
6 Numerical study of the points of intersection
Using the formulas of Subsection 3.4, we look for an equation for $\nu$ as a function of $n$. We recall that for a pair $(x,\nu)$ corresponding to an intersection between $\beta\mapsto\eta(n,\beta)$ and $\beta\mapsto\eta(n+1,\beta)$, we have (3.14), that is
$x=(1-2\nu+n+\frac{1}{2})+\frac{1}{2}\sqrt{(3-4\nu)^{2}+8(1-2\nu)n}\,,$
and (3.5), that is
$(n+1)(n-x)M(\nu,n+1,x)+2x\nu M(\nu+1,n+2,x)=0\,.$
By elimination of $x$, we get the equation
$\Phi(\nu,n)=0\,$ (6.1)
a priori satisfied for $(\nu_{n},n)$ with $n\in\mathbb{N}$ but which could be analyzed for $n\in\mathbb{R}^{+}$.
Using Mathematica, we determined numerically the points of intersection $(\beta_{n},\hat{\eta}_{n})$ of the curves $\beta\mapsto\eta(n,\beta)$ and $\beta\mapsto\eta(n+1,\beta)$ by solving for $(\beta,\eta)$ in the non-linear system consisting of Equations (3.5) and (3.6), with the integer $n$ ranging from 00 to $400$. These points are displayed on Figure 3 as black dots, for $n$ from 0 to $19$. A sample of the results for larger $n$ is presented in Table 1. This computation does not use the Saint-James Formula (3.14). The superscript is used to distinguish these numerical values from those obtained by a second method. We note that the sequence $(\hat{\eta}_{n})$ appears to be increasing.
Table 1: Solutions of the non-linear system
$n$ $\beta_{n}^{(1)}$ ${\hat{\eta}}_{n}^{(1)}$
0 3.847538710016439 0.46188467750410933
1 6.784689992385673 0.490953836999826
2 9.495696565685895 0.5057893193465876
3 12.091164794355297 0.5152514126454681
4 14.613601105384173 0.5219883372745205
5 17.08457097842645 0.5271130898896494
10 28.989490930878333 0.5418512305407657
25 62.88412636538398 0.55750340973811
50 117.3339755112376 0.5663294771262841
100 223.66235051600012 0.5729419029706077
200 432.6371167436942 0.5777978340023635
300 639.5318373766472 0.5799955549150178
400 845.3470994716895 0.5813189732301576
Figure 3: Zoom on the plot of $\beta\mapsto\eta(n,\beta)$ for $n$ from 0 to 20. The black dashed line is the constant $\Theta_{0}$. The black dots mark the points $(\beta_{n}^{(1)},\hat{\eta}_{n}^{(1)})$.
Let us recall that the sequence $\gamma_{n}:=\beta_{n+1}-\beta_{n}$ converges to $2$ and, more precisely, admits an asymptotic expansion of the form
$\gamma_{n}=2+\sum_{j\geq 1}\hat{\gamma}_{j}\,n^{-j/2}+O\left(n^{-\infty}\right)$
(see Equation (4.34)). After applying four times Richardson extrapolation [^8] Assuming that $(y_{n})$ is a sequence such that $y_{n}=\ell+c\,n^{-k/2}+\mathcal{O}\left(n^{-(k+1)/2}\right)$ for some $\ell,\,c\in\mathbb{R}$ and $k\in\mathbb{N}\setminus\{0\}$, the sequence $z_{n}:=\frac{2^{k/2}y_{2n}-y_{n}}{2^{k/2}-1}$ satisfies $z_{n}=\ell+\mathcal{O}\left(n^{-(k+1)/2}\right).$ , we obtain a sequence $(R_{4}\gamma_{n})$ satisfying
$R_{4}\gamma_{n}=2+\mathcal{O}\left(n^{-5/2}\right)$
The new sequence therefore has the same limit as the original, with a much faster convergence, and can be used to check numerically the value of the limit. The results are presented in Table 2.
Table 2: Sequence $\gamma_{n}=\beta_{n+1}-\beta_{n}$
$n$ $\gamma_{n}$ $R_{4}\gamma_{n}$
0 2.9371512823692343
1 2.7110065733002218 1.992309139244719
2 2.595468228669402 1.999385754276766
3 2.5224363110288763 2.0000780253629706
4 2.470969873042275 2.0001506529324526
5 2.4322005977636465 2.000134587773707
6 2.4016457058734133 2.000108698821863
7 2.376764086900799 2.0000863469016394
8 2.355993937178809 2.000068928763249
9 2.338315624735216 2.000055641011863
10 2.3230315388610627 2.00004548367877
24 2.2159762440866757 2.0000068128960815
50 2.1516833824983337
100 2.107942427892027
200 2.076571962353569
300 2.0625877291374763
399 2.0542993010602686
We computed the sequence $(\hat{\eta}_{n}^{(2)})$, for the same values of $n$, by a second method: numerically solving Equation (6.1), which is itself deduced from the Saint-James Formula (3.14). The results are presented in Table 3, with the relative variation from the results of the first method, defined as
$\varepsilon_{n}:=\frac{|{\hat{\eta}}_{n}^{(2)}-{\hat{\eta}}_{n}^{(1)}|}{{\hat{\eta}}_{n}^{(1)}}.$
The close agreement confirms the Saint-James Formula and suggests that our numerical computations are rather accurate.
Table 3: Solutions of the implicit equation, with the relative variation from Table 1
$n$ $\eta_{n}^{(2)}$ $\varepsilon_{n}$
0 3.8475387100164355 $10^{-15}$
1 6.784689992385671 $2\times 10^{-15}$
2 9.495696565685904 $10^{-15}$
3 12.091164794355297 $3\times 10^{-15}$
4 14.613601105384173 $3\times 10^{-15}$
5 17.084570978426473 $2\times 10^{-15}$
10 28.989490930878333 $9\times 10^{-16}$
25 62.884126365384006 $2\times 10^{-15}$
50 117.3339755112376 $2\times 10^{-16}$
100 223.66235051600046 $10^{-15}$
200 432.63711674369415 $10^{-15}$
300 639.5318373766468 $10^{-15}$
400 845.3470994716891 $10^{-15}$
Finally, we present the numerical results concerning the left and right derivatives of $\lambda(\beta)$ at $\beta=\beta_{n}$, obtained with Mathematica, in Table 4.
To compare the limits (5.15) and (5.16) with our numerics, we proceed as we did for the sequence $(\gamma_{n})$, applying four times Richardson extrapolation to the numerical sequences. The results are displayed in Table 4.
All these numerical checks appear consistent with the asymptotic results of Sections 4 and 5.
Table 4: Left and right derivatives at the points of intersection
$n$ $\beta_{n}$ $\lambda^{\prime}(n,\beta_{n})$ $\lambda^{\prime}(n+1,\beta_{n})$ $R_{4}\lambda^{\prime}(n,\beta_{n})$ $R_{4}\lambda^{\prime}(n+1,\beta_{n})$
0 3.84754 0.884743 0.144907
1 6.78469 0.880478 0.178114 0.883010 0.298037
2 9.49570 0.879482 0.195254 0.882910 0.297498
3 12.0912 0.879172 0.206280 0.882882 0.297401
4 14.6136 0.879085 0.214181 0.882872 0.297374
5 17.0846 0.879088 0.220223 0.882868 0.297363
6 19.5168 0.879130 0.225048 0.882866 0.297357
7 21.9184 0.879190 0.229024 0.882861 0.297355
8 24.2952 0.879257 0.232376 0.882864 0.297353
9 26.6512 0.879326 0.235256 0.882864 0.297352
10 28.9895 0.879394 0.237767 0.882863 0.297352
25 62.8841 0.880145 0.256706 0.882863 0.297350
50 117.334 0.880738 0.267540
100 223.662 0.881254 0.275737
200 432.637 0.881671 0.281801
300 639.532 0.881870 0.284558
400 845.347 0.881993 0.286222